In a scale drawing of a house, the living room is 1in. long. The actual length is 24 ft. On the drawing, a hallway is 1/4 in. wide . What is the actual width ?
I had trouble trying to solve this problem , please help . Thanks A Lot ! :)

Answer by stanbon(75887)

In a scale drawing of a house, the living room is 1in. long. The actual length is 24 ft. On the drawing, a hallway is 1/4 in. wide . What is the actual width ?
----
Use a proportion:
ft/in = ft/in
-------------------
x/(1/4) = 24/1
Multiply both sides by 1/4 to get:
x = (1/4)24 = 6 ft.
=======================
cheers,
Stan H.